Please explain!
Chapter 02, Problem 85 Review Int a velocity of 17.0 m/s due north, when a car zooms by at a constant velocity of 41.0 m/s due north. After a reaction time 0.800 s the policeman begins to pursue the speeder with an acceleration of 3.00 m/s 2. Including the reaction time, how long does it take for the police car to catch up with the speeder? ingWare 2.2 in preparation for this problem. A police car is traveling at- VP Speeder VOP -Police car begins
Answer #1

During the reaction time , the police travels = 17*0.8 = 13.6 m

During this time the other car travels = 41*0.8 = 32.8 m

So, relative displacement, do = -(32.8 - 13.6) = -19.2 m

initial relative velocity of the two cars, u = 41 - 17 = 24 m/s

relative acceleration , a = -3 m/s2

Using the equation, s = ut + 0.5*at^2

-19.2 = 24*t + 0.5*(-3)*t^2

So, t = 16.8 s

So, total time taken = 16.8 + 0.8 = 17.6 s <-------- answer